OpenStudy (anonymous):

Please help me solve Rob is 2 more than 3 times teds age. Four years ago rob was 5 times teds age. How old is rob?

OpenStudy (australopithecus):

2 + 3(x) = y 5(y) - 4 = x where x = Rob's age y = Ted's age Use system of equations to solve

OpenStudy (anonymous):

How do you solv this thought

OpenStudy (tkhunny):

Maybe 5(y-4) = x-4? The question does read a little ambiguously.

OpenStudy (australopithecus):

hmm yeah I might have gotten the last equation wrong, but it seems fine to me, because age increases in a linear fashion, so it shouldn't matter if I just subtract 4 from his age

OpenStudy (australopithecus):

To solve a system of equations sub one equation into the other equation using: 2 + 3(x) = y so, 5(y) - 4 = x becomes 5(2 + 3(x)) - 4 = x now solve for x once you have x sub that into one of the two original equations and solve for y

OpenStudy (anonymous):

Is the answer 26

OpenStudy (australopithecus):

Just follow my instructions and you will arrive at the right answer

OpenStudy (anonymous):

Ok